# Ruelle-Pollicott resonances of diffusive U(1)-invariant qubit circuits

### v1

@@ V Conclusion
We have studied the (leading) Ruelle-Pollicott resonances in 3-site translationally invariant qubit circuits with conserved magnetization, which correspond to the (leading) eigenvalues of the truncated operator propagator in the limit of no truncation. Our main result is that the quasi-momentum $k$ dependence of the leading Ruelle-Pollicott resonance $\lambda_{1}$ can be used to extract the transport properties. Specifically, for diffusive systems, we have $\left\lvert\lambda_{1}(k)\right\rvert={\rm e}^{-Dk^{2}}$ for small $k$, allowing us to numerically determine the spin diffusion constant $D$. For large $k$, the leading resonance is not related to transport and governs the decay of the respective correlation functions. We also argue for the existence of a continuum of eigenvalues below the leading diffusive resonance, $\left\lvert\lambda\right\rvert<{\rm e}^{-Dk^{2}}$, which controls non-exponentially decaying correlation functions, for instance, due to power-law hydrodynamic tails.
@@
Even though our analysis was limited to a specific qubit circuit model, we expect our conclusions about the Ruelle-Pollicott resonances to be generic for systems with exactly one U(1) conserved quantity. In particular, the final expression for matrix elements of the truncated propagator in Eq. (10) [or, in case of larger operator spreading, Eq. (28)] hold for any quantum circuit, for example, with a different arrangement of gates or higher local dimension. An interesting open question is generalizing the truncated propagator to Hamiltonian systems.
@@
Another compelling direction is benchmarking the truncated propagator method for extracting the diffusion constant against other standard methods. While we have shown that the Ruelle-Pollicott method produces the same results as a tensor-network simulation and the Green-Kubo formula, it is not clear whether and when one method might be preferable to others. Additionally, the leading Ruelle-Pollicott resonance can hold information about non-diffusive transport, i.e., ballistic, sub- or super-diffusive. However, the exact dependence of the leading resonance on $k$ depends on the asymptotic (space and time) behavior of the correlation function of the local density of the conserved quantity and might be complicated. This can hinder the effectiveness of identifying the transport type and extracting model-dependent constants. Especially interesting are integrable systems with an extensive number of conserved quantities, where we expect an extensive number of RP continuums governing transport.
@@
A focus of future work can also be to better understand the conjecture about subleading Ruelle-Pollicott continuums. A promising direction seems to be developing a different truncation scheme that separates asymptotically differently behaving observables, i.e., those which are governed by hydrodynamics and those which are not. This can also shed additional light on the stretched-exponential decay recently predicted for correlation functions of local non-hydrodynamic observables in 1d systems McCulloch et al. (2025). A particularly interesting physical question in this context is also the asymptotic behavior of the extensive current correlation function. We have shown that it is governed by the corrections to the diffusive correlation function of local magnetization recently argued to be universal for systems with exactly one U(1) conserved quantity Michailidis et al. (2024).
@@
Acknowledgements. U.D. would like to thank Rustem Sharipov, Matija Koterle, Lenart Zadnik, and Tomaž Prosen for insightful discussions. We also acknowledge support by Grants No. J1-4385 and No. P1-0402 from the Slovenian Research Agency (ARIS).
@@ Appendix A Details about the quasi-momentum-dependent propagator
In this appendix, we more carefully define the mathematical structure of the space of extensive observables briefly explained in Sec. II.1 and use it to derive the matrix elements of the truncated propagator, Eq. (10).
@@ A.1 Space of extensive observables
As explained in words in Sec. II.1, the space of extensive observables of support $r$ is spanned by
@@
where $B_{k}^{(m,b)}$ is defined in Eq. (8) and $\mathcal{P}^{(r)}$ in Eq. (6). $\mathbb{P}_{k}^{(r)}$ is an orthonormal basis w.r.t. the extensive Hilbert-Schmidt inner product, $\left\langle\!\left\langle A,B\right\rangle\!\right\rangle=\frac{s}{N2^{N}}\mathrm{tr}\,A^{\dagger}B$, defined in Sec. II.1. This can be deduced from a direct calculation,
@@
for $b,b^{\prime}\in\text{Span}(\mathcal{P}^{(r)})$. Here the local Hilbert-Schmidt inner product is used, $\left\langle a,b\right\rangle=\frac{1}{2^{N}}\mathrm{tr}\,a^{\dagger}b$, defined already in Sec. II.1. This is a consequence of the fact that $b,b^{\prime}$ act with a traceless operator on the first site. For any $j\neq j^{\prime}$ we thus get a traceless operator acting on either site $j$ or $j^{\prime}$, rendering the whole inner product zero. If additionally $b,b^{\prime}\in\mathcal{P}^{(r)}$, we have $\left\langle b,b^{\prime}\right\rangle=\delta_{b,b^{\prime}}$.
@@ A.2 Propagator of extensive observables
In many cases, the sum in Eq. (28) has only a few non-zero terms. The reason for that is similar to the one employed in showing Eq. (27): since $b,b^{\prime}$ act with a traceless on the first site, their translations must overlap on it in order to give a non-zero inner product. In addition to that, Eq. (28) contains propagation $\mathcal{U}\left(\mathcal{S}^{m^{\prime}}(b^{\prime})\right)$, which can change the first site on which $b^{\prime}$ acts non-trivially.
As mentioned already at the beginning of Sec. II, in quantum circuits, the support of observables can increase by at most a finite amount in one period, i.e., the light cone is exact. In general, the spreading of operators depends on the geometry, they can even spread to all sites in one time period, e.g., in staircase circuits. Additionally, it depends on the site at which they act, e.g., a two site operator in a brickwall circuit shown in Fig. 2 spreads by $1$ site to the left and right if it begins at an odd site and by $2$ sites of the left and right if it begins at an even site (the case depicted in the figure). Furthermore, it can also be different to the left and to the right, e.g., in the circuit considered in the bulk of this paper, the operator shown in Fig. 3 spreads by $6$ sites to the left and $4$ sites to the right. To simplify the discussion, we define $\delta r$ to be the maximum spreading for any site and any direction (left/right) in a considered circuit. One can easily see that if $\delta r\leq s$, only three terms in Eq. (28) can be non-zero, and we can simplify it to
$\left[\mathcal{U}_{k}\right]_{(m,b),(m^{\prime},b^{\prime})}=\sum_{j=-1}^{1}{\rm e}^{{\rm i}kj}\left\langle\mathcal{S}^{sj+m}(b),\mathcal{U}\mathcal{S}^{m^{\prime}}(b^{\prime})\right\rangle.$ (29)
This is precisely Eq. (10) we set out to derive. An analogous form was written down in particular cases in Refs Prosen (2007); Žnidarič (2024); Žnidaric et al. (2024). The sum has an interpretation of “realigning” propagated densities back into the basis, where the mapping between densities and extensive observables is one-to-one. For details about this interpretation see Appendix A in Ref. Sharipov et al. (2025).
Eq. (29) applies in many circuits studied in the literature, such as the brickwall circuits, where $\delta r=s=2$, see Fig. 2. It also applies in the whole range of canonical geometries introduced in Ref. Duh and Žnidarič (2024). In the circuits considered in the bulk of this paper (see Fig. 3), the spreading is faster, though; there $\delta r=6$ and $s=3$ and therefore more than $3$ terms in Eq. (28) can be non-zero. One can, however, use a trick described in the following section.
@@ A.3 Space-time symmetries
While matrix elements of the truncated propagator for arbitrary spreading $\delta r$ can be evaluated by Eq. (28), the expressions can sometimes be simplified by exploiting a space-time symmetry. Namely, in both the brickwall circuit defined in Eq. (12) and the 3-site generalization of the brickwall defined in Eq. (14), one can notice that successive layers (i.e., $U_{\mathrm{odd/even}}$ or $U_{1/2/3}$) are translations of the previous layers. This can be interpreted as a space-time symmetry; translations in space are equivalent to translations by a fraction of a period in time. In equation, valid for both of the mentioned cases,
@@
where $S$ is the 1-site translation operator to the left, $S^{\dagger}(a\otimes\mathbbm{1})S=\mathbbm{1}\otimes a$, $s$ is the number of sites of translation invariance (and, crucially, also the number of layers), and $U(t_{1},t_{2})$ denotes the propagator from time $t_{1}$ to $t_{2}$. The time is defined to be propagated by $1/s$ after the application of each layer, in particular the Floquet propagator is $U\equiv U(0,1)$.
Space-time symmetries and their implications are discussed in greater detail Ref. Duh and Žnidarič (2024). An important result for a large class of circuits, including both considered in this paper, is that their propagator can be written in the following way
$U=S^{-s}(S\tilde{U})^{s},$ (31)
where $\tilde{U}$ is a Floquet propagator of a circuit of some other (preferably simpler) geometry. In both of the considered examples, it is the propagator for 1 layer, $\tilde{U}=U_{\mathrm{odd}}$ for the brickwall (12) and $\tilde{U}=U_{1}$ for its 3-site generalization (14).
Since $S^{-s}$ is just a phase in the quasi-momentum eigenbasis, one can redefine $S\tilde{U}$ to be the new equivalent Floquet propagator (i.e., the propagator for one period). In particular, for the 3-site generalization of the brickwall, this reduces the spreading in one period from $\delta r=6$ to $\delta r=2<s=3$, allowing us to use the simpler expression for the matrix elements of the truncated propagator, Eq. (10), and allowing for bigger numerically reachable support (it requires working with operators supported on at most $r+2\delta r=r+4$ sites, instead of $r+12$), see Appendix B.1 for details about the numerics. Since long-time behavior does not depend on the redefinition of the Floquet propagator, the extracted RP resonances must be the same. One must merely account that the new Floquet propagator $S\tilde{U}$ propagates only for time $1/s$ and thus take the $s$-th power of the extracted resonances. Additionally, if one also studies the phase of RP resonances, they must be rotated in each $k$ sector by the appropriate phase stemming from $S^{-s}$ in Eq. (31).
@@ B.1 Spectrum of the truncated propagator
RP resonances are extracted numerically by diagonalizing $\mathcal{U}_{k}^{(r)}$, which can be represented as a finite matrix with matrix elements explicitly given by Eq. (10). Support $r$ basis has $\frac{3s}{4}4^{r}$ elements [see Eq. (25)], thus resulting in a $\frac{3s}{4}4^{r}\times\frac{3s}{4}4^{r}$ matrix. Importantly, the evaluation of matrix elements is performed on propagated observables, which have support $r+2\delta r$ and are, therefore, numerically represented as $\frac{3s}{4}4^{(r+2\delta r)}$ vectors.
Exact diagonalization can be used for smaller supports, up to around $r=7$ for our case of $s=3$. At larger supports, we are constrained by memory and are forced to use iterative methods, which target only a finite number of eigenvalues. Since we are typically interested only in the first few leading eigenvalues, this can be done efficiently. We must only evaluate Eq. (10) on the fly, i.e., by directly acting on a given observable. The calculations are again done on propagated observables with support $r+2\delta r$, which means that the maximum reachable support is constrained also by the amount of spreading. In the considered cases, this approach gives us access to support around $r=13$. We use Arnoldi iteration implemented in ARPACK Lehoucq et al. (1998).
The truncated propagator is generically a non-normal matrix (i.e., it does not commute with its Hermitian adjoint) and its eigenvectors are, therefore, not orthogonal. This can lead to numerical difficulties and subtleties when interpreting its spectrum Trefethen and Embree (2005). It was already observed in Ref. Prosen (2002) that RP resonance eigenvectors are singular objects, however, to what extent non-normality causes other problems when extracting RP resonances remains to be seen.
@@ B.2 Correlation functions
where the sum is taken over some normalized basis of the Hilbert space, and $N$ is the number of sites in the circuit. For clarity, we assumed traceless $A$, although the following discussion simply generalizes to any $A$. Eq. (32) can be numerically evaluated by definition up to $N=18$. In bigger systems, we are constrained by memory and must use the typicality approach. Namely, fluctuations of $C_{AA}(t)$ in a single random state are of the order of $\sim 1/2^{N/2}$. Therefore, the trace in Eq. (32) can be evaluated only in 1 random state, resulting in error on the order of $\sim 3\cdot 10^{-5}$ at $N=30$. Crucially, both $A$ and $U$ must be sufficiently local for this approach and their action on a state must be implemented efficiently. That is, the whole $2^{N}\times 2^{N}$ matrices must not be saved in memory, their action on a state must be evaluated on the fly [e.g., by Eqs (5), (12) and (14)].
@@ B.3 Domain wall quench
To determine the diffusion constant independently of RP resonances, we simulate unitary evolution starting with a weakly polarized domain wall state (see Sec. III.1). In order to probe the behavior in the thermodynamic limit, it is crucial that the numerical simulation only goes up to times, where the boundary has not yet meaningfully affected the dynamics. In diffusive systems, this time can be approximated by the expected spreading of the diffusive front $t\sim x_{\mathrm{max}}^{2}/D$. To reach big enough times, we use TEBD Vidal (2003, 2004). The diffusion constant is expressed from the transferred magnetization in state defined in Eq. (18). According to Eq. (19), this involves evaluating
@@
There are multiple ways to numerically calculate this, for example one can write either $\rho$ or $\sigma_{z}^{j}$ as a matrix product operator (MPO) and evolve them, or appropriately sample and evolve pure states. For small $\mu$, evolving $\rho$ turns out to have the slowest growth of Schmidt coefficients and is therefore the most efficient to simulate. This is compatible with previous findings in similar systems, for details see Ref. Ljubotina et al. (2017). We used this method in the main part of the paper and additionally set the maximum bond dimension to $\chi=256$, discarding other components. Additionally, we checked the results indeed converged by considering different $\chi$ and estimated the error in the diffusion constant $D$ by fitting the prediction in Eq. (19) to different time windows.
Refer to captionRefer to caption(a)(a.i)       Refer to caption (b)
Figure 10: Extraction of the leading RP resonance by a weak-dephasing Lindbladian [see Eq. (35)], an analogous figure to Fig. 2. (a) The leading eigenvalue for the open brickwall quantum circuit with a Haar-random gate $V$. Dashed and dotted lines denote extrapolation to zero noise; the dashed line is a quadratic fit, while the dotted are a linear fit. The inset (a.i) shows the convergence of the second breaking extrapolation (dotted lines) compared to the first breaking extrapolation (dashed lines) and the RP resonance estimated from $r=12$ truncated propagator (the biggest gap is at $k=0$). (b) Infinite temperature autocorrelation function of extensive $k=0$ observables, defined in Eq. (4). The local densities are either $\sigma^{z}$ (i.e., $z$ magnetization) or chosen randomly according to the Gaussian unitary ensemble Mehta (2004). The dashed and dotted lines denote the decays $\propto\left\lvert\lambda\right\rvert^{t}$ governed by the first and second breaking predictions, respectively. The numerics are done in a finite circuit with $N=30$ sites, the $\sigma_{z}$ correlation function is multiplied by $9$ for better presentation.

### v2

@@ 5 Conclusion
We have studied the (leading) Ruelle-Pollicott resonances in translationally invariant qubit circuits with 3-site magnetization-conserving gates. The resonances were obtained from the (leading) eigenvalues of the truncated operator propagator in the limit of no truncation. Our main result is that the quasi-momentum dependence of the leading Ruelle-Pollicott resonance $\lambda_{1}$ can be used to extract the transport properties. Specifically, for diffusive systems, we have $\left\lvert\lambda_{1}(k)\right\rvert={\rm e}^{-Dk^{2}}$ for small quasi-momenta $k$, allowing us to numerically determine the spin diffusion constant $D$. For large $k$, the leading resonance is not related to transport and governs the decay of respective correlation functions. We also argue for the existence of a continuum of eigenvalues below the leading diffusive resonance, $\left\lvert\lambda\right\rvert<{\rm e}^{-Dk^{2}}$, which controls non-exponentially decaying correlation functions, for instance, due to power-law hydrodynamic tails.
@@
Even though our analysis was limited to a specific qubit circuit model, we expect our conclusions about the Ruelle-Pollicott resonances to be generic for systems with exactly one U(1) conserved quantity. In particular, the final expression for the matrix elements of the truncated propagator in Eq. (10) [or, in the case of larger operator spreading, Eq. (28)] hold for any quantum circuit, for example, with a different arrangement of gates or higher local dimension. An interesting open question is generalizing the truncated propagator to Hamiltonian systems.
@@
Another compelling direction is benchmarking the truncated propagator method for extracting the diffusion constant against other standard methods. While we have shown that the Ruelle-Pollicott method produces the same results as a tensor-network simulation and the Green-Kubo formula, it is not clear whether and when one method might be preferable to others [^10] While a detailed analysis would be nice, it seems that the RP approach has some advantages compared to the Lindblad-based approaches with dissipation targeting long Pauli strings, like the DAOE [38]. Namely, we work directly in the infinite-size limit and do the truncation instantly and explicitly. For instance, for a particular system in Fig. 7 of Ref. [51] the RP approach gives the diffusion constant that is within less than $1\%$ of what is believed to be the correct value, while the DAOE is about $3\%$ off (and the Krylov method about $5\%$ off).. Additionally, the leading Ruelle-Pollicott resonance can hold information about non-diffusive transport, i.e., ballistic, sub- or super-diffusive. However, the exact dependence of the leading resonance on $k$ depends on the asymptotic (space and time) behavior of the correlation function of the local density of the conserved quantity (i.e., the scaling function) and might be complicated. This can hinder the effectiveness of identifying the transport type and extracting model-dependent constants. Especially interesting are integrable systems with an extensive number of conserved quantities, where we expect an extensive number of RP continuums governing transport.
@@
A focus of future work can also be to better understand the conjecture about subleading Ruelle-Pollicott continuums. A promising direction seems to be developing a different truncation scheme that separates asymptotically differently behaving observables, i.e., those which are governed by hydrodynamics and those which are not. This can also shed additional light on the stretched-exponential decay recently predicted for correlation functions of local non-hydrodynamic observables in 1d systems [26]. A particularly interesting physical question in this context is also the asymptotic behavior of the extensive current correlation function. We have shown that it is governed by the corrections to the diffusive correlation function of local magnetization recently argued to be universal for systems with exactly one U(1) conserved quantity [28].
Acknowledgements
@@
U.D. would like to thank Rustem Sharipov, Matija Koterle, Lenart Zadnik, and Tomaž Prosen for insightful discussions. We also acknowledge support by Grants No. J1-4385 and No. P1-0402 from the Slovenian Research Agency (ARIS).
@@ Appendix A Details about the quasi-momentum-dependent propagator
In this appendix, we more carefully define the mathematical structure of the space of extensive observables briefly explained in Sec. 2.1 and use it to derive the matrix elements of the truncated propagator, Eq. (10).
@@ A.1 Space of extensive observables
As explained in words in Sec. 2.1, the space of extensive observables of support $r$ is spanned by
@@
where $B_{k}^{(m,b)}$ is defined in Eq. (8) and $\mathcal{P}^{(r)}$ in Eq. (6). $\mathbb{P}_{k}^{(r)}$ is an orthonormal basis w.r.t. the extensive Hilbert-Schmidt inner product, $\left\langle\!\left\langle A,B\right\rangle\!\right\rangle=\frac{s}{N2^{N}}\mathrm{tr}\,A^{\dagger}B$, defined in Sec. 2.1. This can be deduced from a direct calculation,
@@
for $b,b^{\prime}\in\text{Span}(\mathcal{P}^{(r)})$. Here the local Hilbert-Schmidt inner product is used, $\left\langle a,b\right\rangle=\frac{1}{2^{N}}\mathrm{tr}\,a^{\dagger}b$, defined already in Sec. 2.1. This is a consequence of the fact that $b,b^{\prime}$ act with a traceless operator on the first site. For any $j\neq j^{\prime}$ we thus get a traceless operator acting on either site $j$ or $j^{\prime}$, rendering the whole inner product zero. If additionally $b,b^{\prime}\in\mathcal{P}^{(r)}$, we have $\left\langle b,b^{\prime}\right\rangle=\delta_{b,b^{\prime}}$.
@@ A.2 Propagator of extensive observables
In many cases, the sum in Eq. (28) has only a few non-zero terms. The reason for that is similar to the one employed in showing Eq. (27): since $b,b^{\prime}$ act with a traceless Pauli operator on the first site, their translations must overlap on it in order to give a non-zero inner product. In addition to that, Eq. (28) contains propagation $\mathcal{U}\left(\mathcal{S}^{m^{\prime}}(b^{\prime})\right)$, which can change the first site on which $b^{\prime}$ acts non-trivially.
As mentioned already at the beginning of Sec. 2, in quantum circuits, the support of observables can increase by at most a finite amount in one period, i.e., the light cone is sharp. In general, the spreading of operators depends on the geometry, they can even spread to all sites in one time period, e.g., in staircase circuits. Additionally, it depends on the site at which they act, e.g., a two site operator in a brickwall circuit shown in Fig. 2 spreads by $1$ site to the left and right if it begins at an odd site and by $2$ sites to the left and right if it begins at an even site (the case depicted in the figure). Furthermore, it can also be different to the left and to the right, e.g., in the circuit considered in the bulk of this paper, the operator shown in Fig. 3 spreads by $6$ sites to the left and $4$ sites to the right. To simplify the discussion, we define $\delta r$ to be the maximum spreading for any site and any direction (left/right) in a considered circuit. One can easily see that if $\delta r\leq s$, only three terms in Eq. (28) can be non-zero, and we can simplify it to
$\left[\mathcal{U}_{k}\right]_{(m,b),(m^{\prime},b^{\prime})}=\sum_{j=-1}^{1}{\rm e}^{{\rm i}kj}\left\langle\mathcal{S}^{sj+m}(b),\mathcal{U}\left(\mathcal{S}^{m^{\prime}}(b^{\prime})\right)\right\rangle.$ (29)
This is precisely Eq. (10) we set out to derive. An analogous form was written down in particular cases in Refs. [37, 53, 52]. The sum has an interpretation of “realigning” propagated densities back into the basis, where the mapping between densities and extensive observables is one-to-one. For details about this interpretation see Appendix A in Ref. [41].
Eq. (29) applies to many circuits studied in the literature, such as the brickwall circuits, where $\delta r=s=2$, see Fig. 2. It also applies to the whole set of canonical geometries with 2-site nearest-neighbor gates introduced in Ref. [6]. In the 3-site gate circuits considered in the bulk of this paper (see Fig. 3), the spreading is faster, though; there $\delta r=6$ and $s=3$ and therefore more than $3$ terms in Eq. (28) can be non-zero. One can, however, use a trick described in the following section.
@@ A.3 Space-time symmetries
While matrix elements of the truncated propagator for arbitrary spreading $\delta r$ can be evaluated by Eq. (28), the expressions can sometimes be simplified by exploiting a space-time symmetry. Namely, in both the brickwall circuit defined in Eq. (12) and the 3-site generalization of the brickwall defined in Eq. (14), one can notice that successive layers (i.e., $U_{\mathrm{odd/even}}$ or $U_{1/2/3}$) are translations of the previous layer. This can be interpreted as a space-time symmetry; translations in space are equivalent to translations by a fraction of a period in time. In equation, valid for both of the mentioned cases,
@@
where $S$ is the 1-site translation operator to the left, $S^{\dagger}(a\otimes\mathbbm{1})S=\mathbbm{1}\otimes a=\mathcal{S}(a\otimes\mathbbm{1})$, $s$ is the number of sites of translation invariance (and, crucially, also the number of layers), and $U(t_{1},t_{2})$ denotes the propagator from time $t_{1}$ to $t_{2}$. Time is defined to be propagated by $1/s$ after the application of each layer, in particular, the Floquet propagator is $U\equiv U(0,1)$.
Space-time symmetries and their implications are discussed in greater detail in Ref. [6]. An important result for a large class of circuits is that their propagator can be written in the following way
$U=S^{-s}\left(S\tilde{U}\right)^{s},$ (31)
where $\tilde{U}$ is the Floquet propagator of a circuit of some other (preferably simpler) geometry. In both of the considered examples, $\tilde{U}$ is the propagator of the 1st layer, $\tilde{U}=U_{\mathrm{odd}}$ for the brickwall (12) and $\tilde{U}=U_{1}$ for its 3-site generalization (14).
Since $S^{-s}$ is just a phase in the quasi-momentum eigenbasis, one can redefine $S\tilde{U}$ to be the new equivalent Floquet propagator (i.e., the propagator for “one period”). In particular, for the 3-site generalization of the brickwall, this reduces the spreading in one period from $\delta r=6$ to $\delta r=2<s=3$, allowing us to use the simpler expression for the matrix elements of the truncated propagator, Eq. (10), and allowing for bigger numerically reachable supports (it requires working with operators supported on at most $r+2\delta r=r+4$ sites, instead of $r+12$), see Appendix B.1 for details about the numerics. Since the long-time behavior is not affected by the redefinition of the Floquet propagator, the extracted RP resonances must be the same. One must merely account for the fact that the new Floquet propagator $S\tilde{U}$ propagates only for time $1/s$ and thus take the $s$-th power of the extracted resonances. Additionally, if one would study the phase of RP resonances, they would have to be rotated in each $k$ sector by the appropriate phase stemming from $S^{-s}$ in Eq. (31).
@@ B.1 Spectrum of the truncated propagator
RP resonances are extracted numerically from $\mathcal{U}_{k}^{(r)}$, which can be represented as a finite matrix with matrix elements explicitly given by Eq. (10). Support $r$ basis has $\frac{3s}{4}4^{r}$ elements [see Eq. (25)], thus resulting in a $\frac{3s}{4}4^{r}\times\frac{3s}{4}4^{r}$ matrix. Importantly, the evaluation of matrix elements is performed by propagating observables with support $r+2\delta r$ and are, therefore, numerically represented as $\frac{3s}{4}4^{(r+2\delta r)}$ component vectors.
Exact diagonalization can be used for smaller supports, up to around $r=7$ for our case of $s=3$. At larger supports, we are constrained by memory and are forced to use iterative methods, which target only a finite number of eigenvalues. Since we are typically interested only in the first few leading eigenvalues, this can be done efficiently. We must only evaluate Eq. (10) on the fly, i.e., by directly acting on a given observable. The calculations are again done on observables with at most support $r+2\delta r$, which means that the maximum reachable support is constrained also by the amount of spreading. In the considered cases, this approach gives us access to support around $r=13$. We use Arnoldi iteration implemented in ARPACK [20].
The truncated propagator is generically a non-normal matrix (i.e., it does not commute with its Hermitian adjoint) and its eigenvectors are, therefore, not orthogonal. This can lead to numerical difficulties and subtleties when interpreting its spectrum [43]. It was already observed in Ref. [35] that RP resonance eigenvectors are singular objects, however, to what extent non-normality causes other problems when extracting RP resonances remains to be seen.
@@ B.2 Correlation functions
where the sum is taken over some normalized basis of the Hilbert space, and $N$ is the number of sites in the circuit. For clarity, we assumed traceless $A$, although the following discussion straightforwardly generalizes to any $A$. Eq. (32) can be numerically evaluated by definition up to $N=18$. In bigger systems, we are constrained by memory and must use the typicality approach. Namely, fluctuations of $C_{AA}(t)$ in a single random state are of the order of $\sim 1/2^{N/2}$. Therefore, the trace in Eq. (32) can be approximated by only 1 random state, resulting in error on the order of $10^{-5}$ at $N=33$. Crucially, both $A$ and $U$ must be sufficiently local for this approach and their action on a state must be implemented efficiently. That is, the full $2^{N}\times 2^{N}$ matrices must never be saved in memory, only their action on a state must be evaluated on the fly [e.g., by Eqs (5), (12) and (14)].
@@ B.3 Domain wall quench
To determine the diffusion constant independently of RP resonances, we simulate unitary evolution starting with a weakly polarized domain wall state (see Sec. 3.1). In order to probe the behavior in the thermodynamic limit, it is crucial that the numerical simulation only goes up to times, where the boundary has not yet meaningfully affected the dynamics. In diffusive systems, this time can be approximated by the expected spreading of the diffusive front $t\sim x_{\mathrm{max}}^{2}/D$. To reach big enough times, we use TEBD [44, 45]. The diffusion constant is expressed from the transferred magnetization in state defined in Eq. (18). According to Eq. (19), this involves evaluating
@@
There are multiple ways to numerically calculate this, for example one can write either $\rho$ or $\sigma_{z}^{j}$ as a matrix product operator (MPO) and evolve them, or appropriately sample and evolve pure states. For small $\mu$, evolving $\rho$ turns out to have the slowest growth of Schmidt coefficients and is therefore the most efficient to simulate. This is compatible with previous findings in similar systems, for details see Ref. [21]. We used this method in the main part of the paper setting the maximum bond dimension to $\chi=256$. Additionally, we checked the results indeed converged by considering different $\chi$ and estimated the error in the diffusion constant $D$ by fitting the prediction in Eq. (19) to different time windows.